A cosmetics company claims that teenage ladies consume an average of P 150 for cosmetic products. In order to verify this claim, a researcher conducted a survey to 40 teenage ladies and found that their mean monthly cosmetic products is P 146 with a standard deviation of P12. Using a=0.05, test whether the companies claim a realistic or not?
1
Expert's answer
2021-07-19T05:52:28-0400

"n=40 \\

\bar{x}=146 \\

s=12 \\

H_0: \mu=150 \\

H_1: \mu ≠150"

When dealing with a large sample of size n >30 from a population which need not be normal but has finite variance, we can use the central limit theorem to justify using the test for normal populations. Even when "\sigma^2" is unknown we can approximate its value with "s^2" in the computation of the test statistic.

Test-statistic:

"Z= \frac{\bar{x}- \mu}{s / \sqrt{n}} \\

Z= \frac{146-150}{12 / \sqrt{40}}= -2.1 \\

α=0.05"

Two-tailed test

"Z_{crit}=1.96"

The decision rule is: Reject H0 if Z ≤ -1.96 or if Z ≥ 1.96.

"Z= -2.1 < Z_{crit} = -1.96"

Reject H0.

There is enough evidence to conclude that the companies claim is NOT realistic at the 0.05 significance level.